# Dielectric Tensor Prediction for Inorganic Materials Using Latent Information from Preferred Potential

### v1

@@ 2.3.1 Data preparation
We leverage on one of the largest open databases of DFT-calculated crystal structure properties, Materials Project[36], to obtain a wide range of training structures. The dataset (MP v2023.11.1) contains 7277 relaxed materials along with calculated dielectric constants. We clean up the dataset by restricting the element in total dielectric constants $\epsilon_{ij}\in[-10,100]$ for $\forall i,j$ and filtering out structures that contain elements not supported by PFP. Finally there are 6,648 structures retained for model training. The data covers the electronic component $\epsilon_{ij}^{\infty}$, ionic component $\epsilon_{ij}^{0}$, and total component $\epsilon_{ij}$ with a range of $[1.0,96.688]$, $[0.0,90.338]$ and $[1.155,98.889]$, respectively (Fig. 4(a)). Fig. 4(b) exhibits that all seven types of crystal systems are included in our dateset. Among them, the orthorhombic crystal system is the most prevalent, comprising 1,636 instances. In contrast, the triclinic and hexagonal systems are the least represented, with only 440 and 443 instances, respectively. The version of PFP architecture introduced in this work supports 72 elements in total, covering a wide range of elements across the periodic table. Excluding noble gases, He, Ne, Ar, and Kr, 68 elements are covered in the 6,648 available structures (Fig. 4(c)).
The above dataset was partitioned into training, validation, and test subsets in an 0.8:0.1:0.1 ratio for the experimental analysis. To mitigate the impact of random variation, each experiment was independently performed five times with distinct random shuffle data splits. The outcomes are presented as the mean and standard deviation of these five iterations.
Figure 4: (a) $\epsilon^{\infty}-\epsilon^{0}$ map of structures in the dataset, colored by $\epsilon$; (b)Counts of crystal system types in teh dataset; (c) Element counts for atoms in the structures with available dielectric properties on Materials Project. Elements under blue are not supported by PFP. Elements under white are supported by PFP but do not appear in the dataset.
2.3.2 Performance of DTNN
@@
These MAE metrics help assess the accuracy of the model predictions for the dielectric tensor properties. In this work, all models are trained using $\mathrm{MAE_{ten}}$ as the loss function. Due to structural symmetry, the number of independent elements in the dielectric tensor varies according to system type (see Table1). Fig. 5(a) shows that generally the intermediate knowledge from the deeper GCN layer contributes more to the dielectric prediction. When PFP-L4 and PFP-L5 are taken as the feature generator for the child model, lower $\mathrm{MAE_{ten}}$ prediction errors are observed on the test structures. This could be explained by that the shallower layers only capture shorter-range local interactions. Meanwhile, quite close accuracy between $n=4$ and $n=5$ is because only their scalar representations are distinct. Considering the better computational efficiency, PFP-L4 is taken as the feature generator and is utilized for the subsequent study. Fig. 5(b-d) showcases the performance of DTNN on each crystal system for the prediction task of the electronic, ionic and total dielectric constants. Due to structural symmetry, Three models DTNN-$\epsilon^{\infty}$, DTNN-$\epsilon^{0}$ and DTNN-$\epsilon$ are separately trained with their $\mathrm{MAE_{ten}},\mathrm{MAE_{mean-diag}},\mathrm{MAE_{diag}},\mathrm{MAE_{off-diag}}$ evaluated. Dielectric tensors of cubic, hexagonal, orthorhombic, tetragonal and trigonal systems hold 0 on their off-diagonal elements due to geometric symmetry. The zero off-diagonal errors for them demonstrates the our model keeps equivariance for correct prediction of 0 on all their off-diagonal elements. Additionally, the same $\mathrm{MAE_{ten}}$ and $\mathrm{MAE_{mean-diag}}$ of the cubic system further confirm the output equivariance is guaranteed due to its symmetric structure along all three directions in the Euclidean space. The model performs uniformly for different crystal systems according to all $\mathrm{MAE_{ten}}$ in the range of $[0.27,0.57]$ for $\boldsymbol{\epsilon^{\infty}}$, $[1.07,1.93]$ for $\boldsymbol{\epsilon^{0}}$ and $[1.41,2.14]$ for $\boldsymbol{\epsilon}$.
Crystal Number of independent elements Dielectric tensor
Cubic 1 $\begin{bmatrix}\epsilon&0&0\\ 0&\epsilon&0\\ 0&0&\epsilon\end{bmatrix}$
Tetragonal, Triagonal, Hexagonal 2 $\begin{bmatrix}\epsilon_{1}&0&0\\ 0&\epsilon_{1}&0\\ 0&0&\epsilon_{3}\end{bmatrix}$
Orthorhombic 3 $\begin{bmatrix}\epsilon_{1}&0&0\\ 0&\epsilon_{2}&0\\ 0&0&\epsilon_{3}\end{bmatrix}$
Monoclinic 4 $\begin{bmatrix}\epsilon_{11}&0&\epsilon_{13}\\ 0&\epsilon_{2}&0\\ \epsilon_{13}&0&\epsilon_{33}\end{bmatrix}$
Triclinic 6 $\begin{bmatrix}\epsilon_{11}&\epsilon_{12}&\epsilon_{13}\\ \epsilon_{12}&\epsilon_{22}&\epsilon_{23}\\ \epsilon_{13}&\epsilon_{23}&\epsilon_{33}\end{bmatrix}$
Figure 5: Performance of DTNN models on dielectric data from Materials Project. (a) Plot of the mean error (points) with standard deviation (shadow) for 5 runs against the intermediate embedding generated from the PFP-Ln as DTNN inputs. (b) 5-run integrated plot of average $\mathrm{MAE_{ten}}$ for each element in the test set against the number of structures containing that element in the training set. (c-e) Multiple MAE metrics in the dielectric tensor for different crystal systems on the prediction of the electronic, ionic and total task.
2.3.3 Model comparison
@@
There are limited open-source works that discussed the prediction of tensorial properties of materials. The models of PAINN[30], M3GNet[34] and MatTen[35] are chosen for comparison in this work. They use different approaches to produce rank-2 equivariant tensors by incorporation of equivariant atomwise representations, partial derivative with respect to the Cartesian tensor, and spherical harmonics expansion, respectively. PAINN encodes each atom $i$ with an invariant representation $s_{i}$ and an equivariant vector representation $v_{i}$. Incorporating $v_{i}$ has been demonstrated to be effective for structure recognition through ablation studies. The dielectric tensor in PAINN is constructed using a rank-1 tensor decomposition, and the atom positions are incorporated to hold the global structure information:
$\epsilon=\sum_{i=1}^{N}\epsilon_{0}(s_{i})I_{3}+\overrightarrow{\nu}(\overrightarrow{v}_{i})\otimes\overrightarrow{r}_{i}+\overrightarrow{r}_{i}\otimes\overrightarrow{\nu}(\overrightarrow{v}_{i})$ (7)
@@
where $I_{3}$ is an identity matrix with size of three, $\epsilon_{0}$ and $\overrightarrow{\nu}$ are invariant and equivariant nonlinearities, respectively. The M3GNet model is able to generate equivariant $3\times 3$ tensor from the partial derivatives of its scalar output using the lattice matrix information. This scalar output is not directly trained as a target variable, and instead we train the model based on the loss computed from the derivative values of this output. Specifically, we calculate the loss of $\mathrm{MAE}_{ten}$ between these derivative results and labeled targets. After training, these derivative values are utilized for dielectric prediction. In MatTen, the atom feature contains a set of scalars, vectors, and high-rank tensors which are produced by spherical harmonics, namely, a geometric object consisting of a direct sum of irreducible representations of the SO(3) rotation group. Its interaction blocks follow the design of Tensor Field Network[38] and NequIP[29]. The second-order features of all atoms are ultimately aggregated as the prediction of dielectric tensor. To investigate the necessity of including edge-related representations generated by PFP-L4, we trained two models, DTNN and DTNN-simple. DTNN leverages all latent embeddings $(\mathbf{a}_{s}^{n},\mathbf{a}_{v}^{n},\mathbf{a}_{t}^{n},\mathbf{b}_{s}^{n},\mathbf{b}_{v}^{n})$, while DTNN-simple only utilizes $(\mathbf{a}_{s}^{n},\mathbf{a}_{v}^{n},\mathbf{a}_{t}^{n})$ and the layers related to Eq.14 and Eq.17 are removed.
@@
Table 2 provides a comparison of the three aforementioned models with DTNN and DTNN-simple. In the prediction task across different systems, the models are trained on the entire training data, while the test structures are categorized according to crystal systems. It is shown that DTNN performs best in 22 among all 24 tasks in comparison with all models. The aggregation of potential edge information showcases significant improvements in accuracy, as evidenced by DTNN beating DTNN-simple in 18 out of the 24 tasks. It should be noted that DTNN-simple still gives much better accuracy than the other 3 models, outperforming 22 tasks. Only two tasks lost to M3GNet when making ionic and total dielectric tensor predictions on triclinic systems. Hence, the latent edge scalar and vector features provide additional useful geometric information which benefits the dielectric prediction.
Models DTNN DTNN-simple PaiNN M3GNet MatTen
Task Electronic ($\epsilon_{\infty}$)
@@
Task Ionic ($\epsilon_{0}$)
@@
Task Total ($\epsilon$)
@@
Table 2: Performance comparison of models across all crystal systems. Models are trained separately for 3 tasks: the electronic component, ionic component and total dielectric constants. The results are shown by the $\mathrm{MAE_{ten}}$ and standard deviation of 5 runs using different random split. The 5 trained DTNN models will be employed as ensemble models for virtual screening in the next section.
@@ 2.4.1 Potential application
Dielectrics are crucial functional materials for in microelectronic device manufacturing. Materials with high dielectric constants have the potential to enable higher energy density storage, leading to performance enhancement and device miniaturization.[39]. Meanwhile, large band gaps ($E_{g}$) are also desirable as they prevent leakage currents when materials are exposed to large electric fields, especially at nanoscale thicknesses.. Hence, novel dielectric materials with both high dielectric constants $\epsilon$ and large band gaps $E_{g}$ are highly preferred for such applications. The DTNN model has been proven effective predicting dielectric tensors which account for the direction and magnitude of the electric field. To represent the capacity of electric energy storage, dielectric tensors $\boldsymbol{\epsilon}$ can be simplified to scalar quantities by Eq.2. To prepare the candidate set for screeening, we downloaded 14,375 non-metal materials from the MP database. These materials are specifically selected based on the energy above convex hull $E_{\mathrm{hull}}=0$ to estimate their thermodynamic stability, so that only stable materials are included in the candidate set. The candidates have unknown dielectric constant properties, and there is no overlap between this candidate set and our training dataset. We attempt to leverage our models for virtual screening on all these candidates to identify potential materials suitable for dielectric applications in microelectronic device manufacturing. To improve the accuracy, we use the ensemble results from the 5 DTNN models trained in the previous section.
2.4.2 Inversely related $E_{g}$ and $\epsilon$
@@
It is known that there exists inverse relationship between the electronic part of dielectric constants $\epsilon^{\infty}$ and band gaps $E_{g}$. On the other hand, the ionic contribution $\epsilon^{0}$ does not show a clear correlation with $E_{g}$. The theoretical background of this phenomenon is explained in reference[40]. This trend is also observed in the materials present in the training set as shown in Fig. 6(a) and Fig. 6(b). To showcase the capability of our model in capturing such latent quantum mechanism knowledge underlying chemical structures, we employ the DTNN-$\epsilon^{\infty}$ and DTNN-$\epsilon^{0}$ models to predict the two dielectric constant components for the 14,375 candidates, respectively. The density of data points are drawn in contours for structures in the training set (red) and candidate set (blue) using the kernel density estimation. Despite the fact that the model does not learn explicitly from band gap data, the distribution of model-predicted electronic dielectric constants of the candidates and their DFPT-computed band gaps from MP indicates the model’s capacity to comprehend this negative relationship between the two properties through materials’ graph representations. This phenomenon can be explained by the interplay between electronic polarization and electronic excitations. Materials with smaller band gaps allow for easier excitation of electrons to higher energy levels, leading to higher electronic polarizability and larger electronic dielectric constants. Conversely, materials with larger band gaps hinder electron excitation, resulting in lower electronic dielectric constants [41]. In contrast to the case of $\epsilon^{\infty}$, no discernible correlation is observed between DTNN-$\epsilon^{0}$-predicted ionic part and band gaps. From a microscopic perspective, the ionic dielectric constant is influenced by the displacement of cations and anions from their equilibrium positions under the influence of electric fields. The strength of bonds or phonon frequencies is highly sensitive to variations in bond lengths. Consequently, materials with similar compositions can exhibit a wide range of $\epsilon^{0}$ due to differences in bond lengths [42]. The weak constraint of $E_{g}$ to $\epsilon^{0}$ along with the broad value spectrum of $\epsilon^{0}$ present the opportunity for the exploration and discovery of novel dielectric materials characterized by both high $\epsilon$ and $E_{g}$.
2.4.3 Virtual screening & DFPT validation
We conducted the virtual screening by predicting total dielectric constant scalar $\epsilon$ for all structures in the candidate set. Following the previous work[42], $E_{g}\cdot\epsilon$ is assigned as the figure of merit (FOM) to quantify the performance of dielectrics, since $E_{g}$ and $\epsilon$ are approximately proportional to the logarithm of the leakage current density[43]. After prediction, we ranked the structures according to FOM and investigated the compositions of the top-100 materials subset compared to all materials in Fig.6(c). Excluding common non-metallic elements like oxygen and fluorine, we found tantalum (Ta) to be the most frequently occurring metallic element among the top candidate materials, followed closely by lead (Pb), bismuth (Bi), and niobium (Nb). To validate the promising materials identified by DTNN, we selected 12 candidates from the top 100 materials based on computational budget constraints. These candidates were chosen to have a unit cell size with a maximum of 20 atoms to enable efficient computation of their dielectric constants using DFPT. The details of DFPT configurations are available in the Method section.
@@ 2.4.3 Virtual screening & DFPT validation
Fig. 6(d) illustrates the distribution of the figure of merit (FOM) for stable materials in both the training set and candidate set. The thermaldynamically stable material with the highest FOM in the training dataset is Sn2ClF3 (mp-29252, $E_{g}=3.56\mathrm{eV},\epsilon^{\infty}=3.81,\epsilon^{0}=72.77,\epsilon=76.55$). We successfully identified a new material Ba2SmTaO6 (mp-1214622, $E_{g}=3.36\mathrm{eV},\epsilon^{\infty}=4.85,\epsilon^{0}=88.96,\epsilon=93.81$) with structural visualized in Fig. 6(e). It FOM score is higher than any calculated stable structures in our database that are used for model training. Three of the constituent elements present in Ba2SmTaO6 (O, Ta, and Ba) are among the most commonly occurring elements in the compositions of materials within the top 100 candidates. This observation further supports the compatibility of these elements in forming high-performance dielectric materials.

### v2

@@ 2.3.1 Data preparation
We leverage on one of the largest open databases of DFT-calculated crystal structure properties, Materials Project [38], to obtain a wide range of training structures. The dataset (MP v2023.11.1) contains 7277 relaxed materials along with calculated dielectric constants. We clean up the dataset by restricting the element in total dielectric constants $\varepsilon_{ij}\in[-10,100]$ for $\forall i,j$ and filtering out structures that contain elements not supported by PFP. Finally there are 6,648 structures retained for model training. The data covers the electronic component $\varepsilon^{\infty}$, ionic component $\varepsilon^{0}$, and total component $\varepsilon$ with its scalar values in the range of $[1.0,96.688]$, $[0.0,90.338]$ and $[1.155,98.889]$, respectively (Fig. 4(a)). Fig. 4(b) exhibits that all seven types of crystal systems are included in our dateset. Among them, the orthorhombic crystal system is the most prevalent, comprising 1,636 instances. In contrast, the triclinic and hexagonal systems are the least represented, with only 440 and 443 instances, respectively. The version of PFP architecture introduced in this work supports 72 elements in total, covering a wide range of elements across the periodic table. Excluding noble gases, He, Ne, Ar, and Kr, 68 elements are covered in the 6,648 available structures (Fig. 4(c)).
The above dataset was partitioned into training, validation, and test subsets in an 0.8:0.1:0.1 ratio for the experimental analysis. To mitigate the impact of random variation, each experiment was independently performed five times with distinct random data splits. The outcomes are presented as the mean and standard deviation of these five iterations.
Figure 4: (a) $\varepsilon^{\infty}-\varepsilon^{0}$ map of structures in the dataset, colored by $\varepsilon$; (b)Counts of crystal system types in teh dataset; (c) Element counts for atoms in the structures with available dielectric properties on Materials Project. Elements under blue are not supported by PFP. Elements under white are supported by PFP but do not appear in the dataset.
2.3.2 Performance of DTNet
Crystal Number of independent elements Dielectric tensor
Cubic 1 $\begin{bmatrix}\varepsilon&0&0\\ 0&\varepsilon&0\\ 0&0&\varepsilon\end{bmatrix}$
Tetragonal, Triagonal, Hexagonal 2 $\begin{bmatrix}\varepsilon_{1}&0&0\\ 0&\varepsilon_{1}&0\\ 0&0&\varepsilon_{3}\end{bmatrix}$
Orthorhombic 3 $\begin{bmatrix}\varepsilon_{1}&0&0\\ 0&\varepsilon_{2}&0\\ 0&0&\varepsilon_{3}\end{bmatrix}$
Monoclinic 4 $\begin{bmatrix}\varepsilon_{11}&0&\varepsilon_{13}\\ 0&\varepsilon_{2}&0\\ \varepsilon_{13}&0&\varepsilon_{33}\end{bmatrix}$
Triclinic 6 $\begin{bmatrix}\varepsilon_{11}&\varepsilon_{12}&\varepsilon_{13}\\ \varepsilon_{12}&\varepsilon_{22}&\varepsilon_{23}\\ \varepsilon_{13}&\varepsilon_{23}&\varepsilon_{33}\end{bmatrix}$
Figure 5: Performance of DTNet models on dielectric data from Materials Project. (a-b) Plot of the mean error/correlation (points) with standard deviation (shadow) for 5 runs against the intermediate embedding generated from the PFP-Ln as DTNet inputs, showing both transfer learning results and training from scratch results for comparison. (c-e) Multiple MAE metrics in the dielectric tensor for different crystal systems on the prediction of the electronic, ionic and total task.
@@ 2.3.2 Performance of DTNet
These MAE metrics help assess the accuracy of the model predictions for the dielectric tensor properties. In this work, all models are trained using $\mathrm{MAE_{ten}}$ as the loss function. Due to structural symmetry, the number of independent elements in the dielectric tensor varies according to system types (see Table1). Fig. 5(a-b) shows that generally the intermediate knowledge from the deeper GCN layer contributes more to the dielectric prediction. When PFP-L4 and PFP-L5 are taken as the feature generator for the child model, lower $\mathrm{MAE_{ten}}$ prediction errors are observed on the test structures. This could be explained by that the shallower layers only capture shorter-range local interactions. Meanwhile, quite close accuracy between $n=4$ and $n=5$ is because only their scalar representations are distinct. Considering the better computational efficiency, PFP-L4 is taken as the feature generator and is utilized for the subsequent study. The comparison of error metrics between models utilizing transfer learning and those trained from scratch highlights the significance of incorporating compositional and structural encodings from PFP. Specifically, the use of pretrained PFP-L4 resulted in a reduction of 10.0% for RMSE and 14.1% in MAE compared to a PFP-L4 model with the same architecture but trained with all parameters initialized from scratch. Fig. 5(c-e) showcases the performance of DTNet on each crystal system for the prediction task of the electronic, ionic and total dielectric constants. Three models DTNet-$\varepsilon^{\infty}$, DTNet-$\varepsilon^{0}$ and DTNet-$\varepsilon$ are separately trained with their $\mathrm{MAE_{ten}},\mathrm{MAE_{mean-diag}},\mathrm{MAE_{diag}},\mathrm{MAE_{off-diag}}$ evaluated. Dielectric tensors of cubic, hexagonal, orthorhombic, tetragonal and trigonal systems hold 0 on their off-diagonal elements due to geometric symmetry. The zero off-diagonal errors for them demonstrates the our model keeps equivariance for correct prediction of 0 on all their off-diagonal elements. Additionally, the same $\mathrm{MAE_{ten}}$ and $\mathrm{MAE_{mean-diag}}$ of the cubic system further confirm the output equivariance is guaranteed due to its symmetric structure along all three directions in the Euclidean space. The model performs uniformly for different crystal systems according to all $\mathrm{MAE_{ten}}$ in the range of $[0.27,0.57]$ for $\boldsymbol{\varepsilon^{\infty}}$, $[1.07,1.93]$ for $\boldsymbol{\varepsilon^{0}}$ and $[1.41,2.14]$ for $\boldsymbol{\varepsilon}$.
Fig. 6(a-b) illustrates the performance of DTNet by calculating the polycrystalline dielectric constants $\varepsilon$ and tensor eigenvalues $\lambda_{i}$ from the predicted dielectric tensors, with the eigenvalues being crucial for evaluating anisotropy. The ground trues of $\varepsilon$ and $\lambda_{i}$ in the test set fall within the range of [2.129, 98.889] and [1.460, 98.889], respectively. DTNet provides predictions that closely match DFT reference values, with MAE of 2.568 for $\varepsilon$ and 3.740 for $\lambda_{i}$, within a prediction range of 96.760 and 97.429, respectively. To further validate the accuracy of our method, we evaluated its performance on the dielectric task using the Matbench dataset [22], which is a subset of structures from the Materials Project. In Matbench, we assessed our method by calculating refractive index $\eta$ from the dielectric tensors predicted by DTNet.
Fig. 7 compares the mean MAEs with error bars representing standard deviations and the mean RMSEs across 5 runs, as defined by the Matbench pipeline. DTNet surpasses all other methods published on the official Matbench leaderboards for the dielectric task, achieving an average MAE of 0.2371 with a standard deviation of 0.0765, and an RMSE of 1.6830 over the 5-fold test data. In comparison, the top-performing model MODNet [40] reported on Matbanch has an MAE of 0.2711 with a standard deviation of 0.0714 and an RMSE of 1.6832.
Figure 6: Performance of DTNet on the test set of our refined MP dataset for two distinct tasks: (a) polycrystalline dielectric constants and (b) tensor eigenvalues.
Figure 7: Benchmarking our approach against other algorithms on the Matbench dataset. All results are sourced from the official Matbench leaderboards. For the MODNet, coGN, and Finder models, we present only their best reported results.
2.3.3 Model comparison for dielectric tensor prediction
@@
At the time of conducting this work, there were limited open-source works that discussed the prediction of tensorial properties of materials. The models of PaiNN [31], M3GNet [36] and MatTen [37] are chosen for comparison in this work. They use different approaches to produce rank-2 equivariant tensors by incorporation of equivariant atomwise representations, partial derivative with respect to the Cartesian tensor, and spherical harmonics expansion, respectively. PaiNN encodes each atom $i$ with an invariant representation $\boldsymbol{s}_{i}$ and an equivariant vector representation $\boldsymbol{v}_{i}$. Incorporating $\boldsymbol{v}_{i}$ has been demonstrated to be effective for structure recognition through ablation studies. The dielectric tensor in PaiNN is constructed using a rank-1 tensor decomposition, and the atom positions $\boldsymbol{r}_{i}$ are incorporated to hold the global structure information:
$\boldsymbol{\varepsilon}=\sum_{i=1}^{N^{a}}s(\boldsymbol{s}_{i})\boldsymbol{I}_{3}+\nu(\boldsymbol{v}_{i})\otimes\boldsymbol{r}_{i}+\boldsymbol{r}_{i}\otimes\nu(\boldsymbol{v}_{i})$ (7)
@@
where $\boldsymbol{I}_{3}$ is an identity matrix with size of three, $s$ and $\nu$ are invariant and equivariant nonlinearities, respectively. The invariant M3GNet model is able to generate equivariant $3\times 3$ tensor from the partial derivatives of its scalar output using the lattice matrix information. This scalar output is not directly trained as a target variable, and instead we train the model based on the loss computed from the derivative values of this output. Specifically, we calculate the loss of $\mathrm{MAE}_{ten}$ between these derivative results and labeled targets. After training, these derivative values are utilized for dielectric prediction. In MatTen, the atom feature contains a set of high-rank tensors produced by spherical harmonics, namely, a geometric object consisting of a direct sum of irreducible representations of the SO(3) rotation group. Its interaction blocks follow the design of Tensor Field Network [41] and NequIP [30]. The second-order features of all atoms are ultimately aggregated as the prediction of dielectric tensor. To investigate the necessity of including edge-related representations generated by PFP-L4, we trained two models, DTNet and DTNet-simple. DTNet leverages all latent embeddings $(\boldsymbol{a}_{s}^{n},\boldsymbol{a}_{v}^{n},\boldsymbol{a}_{t}^{n},\boldsymbol{b}_{s}^{n},\boldsymbol{b}_{v}^{n})$, while DTNet-simple only utilizes $(\boldsymbol{a}_{s}^{n},\boldsymbol{a}_{v}^{n},\boldsymbol{a}_{t}^{n})$ and the layers related to Eq.14 and Eq.17 are removed.
@@
Table 2 provides a comparison of the three aforementioned models with DTNet and DTNet-simple. In the prediction task across different systems, the models are trained on the entire training data, while the test structures are categorized according to crystal systems. It is shown that DTNet performs best in 22 among all 24 tasks in comparison with all models. The aggregation of potential edge information showcases significant improvements in accuracy, as evidenced by DTNet beating DTNet-simple in 18 out of the 24 tasks. It should be noted that DTNet-simple still gives much better accuracy than the other 3 models, outperforming 22 tasks. Only two tasks lost to M3GNet when making ionic and total dielectric tensor predictions on triclinic systems. Hence, the latent edge scalar and vector features provide additional useful geometric information which benefits the dielectric prediction.
Table 2: Performance comparison of models across all crystal systems. Models are trained separately for 3 tasks: the electronic component, ionic component and total dielectric constants. The results are shown by the $\mathrm{MAE_{ten}}$ and standard deviation of 5 runs using different random split. The 5 trained DTNet models will be employed as ensemble models for virtual screening in the next section.
Models DTNet DTNet-simple PaiNN M3GNet MatTen
Task Electronic ($\varepsilon_{\infty}$)
@@
Task Ionic ($\varepsilon_{0}$)
@@
Task Total ($\varepsilon$)
@@ 2.4.1 Preparation of candidate structures
The DTNet model has been proven effective predicting dielectric tensors which account for the direction and magnitude of the electric field. To prepare the candidate set for screening, we downloaded 14,375 non-metal materials from the MP database. These materials are specifically selected based on the energy above convex hull $E_{\mathrm{hull}}=0$ to estimate their thermodynamic stability, so that only stable materials are included in the candidate set. The candidates have unknown dielectric constant properties, and there is no overlap between this candidate set and our training dataset. We aim to leverage our models for virtual screening on all these candidates to identify potential materials for high-dielectric materials and highly anisotropic materials, incorporating both electronic and ionic contributions. To improve accuracy and assess model uncertainty, we use ensemble results from the five DTNet models trained in the previous section. The expected property $\mu(\chi)$ is estimated by averaging the predictions, while the uncertainty $\sigma(\chi)$ is quantified by the standard deviations. We conducted iterative active learning to identify top candidates through virtual screening. See details in the Method section.
2.4.2 High-dielectric materials
Dielectrics are crucial functional materials for microelectronic device manufacturing. Materials with high dielectric constants have the potential to enable higher energy density storage, leading to performance enhancement and device miniaturization [42]. Meanwhile, large band gaps ($E_{g}$) are also desirable as they prevent leakage currents when materials are exposed to large electric fields, especially at nanoscale thicknesses. Hence, novel dielectric materials with both high dielectric constants $\varepsilon$ and large band gaps $E_{g}$ are highly preferred for such applications. To represent the capacity of electric energy storage, dielectric tensors $\boldsymbol{\varepsilon}$ can be simplified to scalar quantities by Eq.2.
@@ 2.4.2 High-dielectric materials
It is known that there exists inverse relationship between the electronic part of dielectric constants $\varepsilon^{\infty}$ and band gaps $E_{g}$. On the other hand, the ionic contribution $\varepsilon^{0}$ does not show a clear correlation with $E_{g}$. The theoretical background of this phenomenon is explained in reference [43]. This trend is also observed in the materials present in the training set as shown in Fig. 8(a) and Fig. 8(b). To showcase the capability of our model in capturing such latent quantum mechanism knowledge underlying chemical structures, we employ the DTNet-$\varepsilon^{\infty}$ and DTNet-$\varepsilon^{0}$ models to predict the two dielectric constant components for the 14,375 candidates, respectively. The density of data points are drawn in contours for structures in the training set (red) and candidate set (blue) using the kernel density estimation. Despite the fact that the model does not learn explicitly from band gap data, the distribution of model-predicted electronic dielectric constants of the candidates and their DFPT-computed band gaps from MP indicates the model’s capacity to comprehend this negative relationship between the two properties through materials’ graph representations. This phenomenon can be explained by the interplay between electronic polarization and electronic excitations. Materials with smaller band gaps allow for easier excitation of electrons to higher energy levels, leading to higher electronic polarizability and larger electronic dielectric constants. Conversely, materials with larger band gaps hinder electron excitation, resulting in lower electronic dielectric constants  [44]. In contrast to the case of $\varepsilon^{\infty}$, no discernible correlation is observed between DTNet-$\varepsilon^{0}$-predicted ionic part and band gaps. From a microscopic perspective, the ionic dielectric constant is influenced by the displacement of cations and anions from their equilibrium positions under the influence of electric fields. The strength of bonds or phonon frequencies is highly sensitive to variations in bond lengths. Consequently, materials with similar compositions can exhibit a wide range of $\varepsilon^{0}$ due to differences in bond lengths  [45]. The weak constraint of $E_{g}$ to $\varepsilon^{0}$ along with the broad value spectrum of $\varepsilon^{0}$ present the opportunity for the exploration and discovery of novel dielectric materials characterized by both high $\varepsilon$ and $E_{g}$.
We conducted the virtual screening by employing DTNet-$\varepsilon$ to predict total dielectric constant scalar $\varepsilon$ for all structures in the candidate set. Following the previous work [45], $E_{g}\cdot\varepsilon$ is assigned as the figure of merit (FOM) to quantify the performance of dielectrics, since $E_{g}$ and $\varepsilon$ are approximately proportional to the logarithm of the leakage current density [46]. To validate the promising materials identified by DTNet, we conducted a three-round active exploration, with each round proposing 20 top candidates for validation. The details of virtual screening and DFPT configurations are available in the Method section. Finally, totally 35 out of the 60 structures across 3 iterations are successfully converged through first-principle calculations.
@@
Fig. 8(c) illustrates the distribution of the figure of merit (FOM) for stable materials in both the training set and screened candidate set. The thermaldynamically stable material with the highest FOM in the training dataset is Sn2ClF3 (mp-29252, $E_{g}=3.56\mathrm{eV},\varepsilon^{\infty}=3.81,\varepsilon^{0}=72.77,\varepsilon=76.55$). We successfully identified 3 new materials, Cs2Ti(WO4)3 (mp-1226157, $E_{g}=2.83\mathrm{eV},\varepsilon^{\infty}=5.34,\varepsilon^{0}=175.55,\varepsilon=180.89$), RbNbWO6 (mp-1219587, $E_{g}=2.93\mathrm{eV},\varepsilon^{\infty}=5.27,\varepsilon^{0}=150.30,\varepsilon=155.57$) and Ba2SmTaO6 (mp-1214622, $E_{g}=3.36\mathrm{eV},\varepsilon^{\infty}=4.85,\varepsilon^{0}=88.96,\varepsilon=93.81$), with structures visualized in Fig. 8(d). Their FOM scores are higher than any calculated stable structures in our training database.